OpenStudy (anonymous):

\[\int\limits \frac{x}{x ^{2}+6x}dx\]

OpenStudy (anonymous):

Nevermind I should read more carefully, this identity is equal to \[\frac{1}{x+6}\]

OpenStudy (anonymous):

how???

OpenStudy (anonymous):

So you will get \[ \large \int \frac{1}{x+6} dx = \ln |x+6| +C\]

OpenStudy (anonymous):

\[ \frac{x}{x^2+6}=\frac{x}{x(x+6)}=\frac{1}{x+6}\]

OpenStudy (anonymous):

the first one should be x^2+6x

OpenStudy (anonymous):

thanks. :)

OpenStudy (anonymous):

welcome